trentin:

Solve for x: −3|2x + 6| = −12

trentin:

@Vocaloid

Vocaloid:

First divideboth sides by -3

Vocaloid:

|2x + 6| = 4 Then solve for the positive and negative condition

Vocaloid:

2x + 6 = 4 2x + 6 = -4

trentin:

you are go to fast

Vocaloid:

ok I can start over

trentin:

ok

Vocaloid:

Since we have -3|2x+6| we want to try and get the absolute value expression by itself

trentin:

ok

Vocaloid:

In order to get rid of that -3 we divide by -3

trentin:

-3 divide -3

Vocaloid:

-3|2x + 6| divided by -3

Vocaloid:

That leaves us with |2x + 6|

trentin:

did you divided the -12 with the -3

Vocaloid:

Yes

trentin:

4

Vocaloid:

|2x + 6| = 4

Vocaloid:

So |2x + 6| = 4 means: 2x + 6 = 4 2x + 6 = -4

Vocaloid:

Then you would solve for x (you should get two x values)

Vocaloid:

|A| = B Implies that A can be B or -B

trentin:

absolute value can be negative

Vocaloid:

The expression inside the absolute value can be negative

Vocaloid:

This is ok: |-A| = 4 This has no solution: |A| = -4

trentin:

ok

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid So |2x + 6| = 4 means: 2x + 6 = 4 2x + 6 = -4 \(\color{#0cbb34}{\text{End of Quote}}\)

Vocaloid:

Can you try solving for x now?

trentin:

yes

trentin:

x=−5,−1

Vocaloid:

Food that's your answer

Vocaloid:

*good not food